3f(x)= f(x+1)+ (5/f(x+2))
then lim x tends to infinity f(x)=?

Grade:11

1 Answers

hardik joshi
11 years ago

HERE, IS

LIM X TENDS TO INFINITIVE

SO,THAT

INFINITIVE+1=INFINITIVE

INFINITIVE+2=INFINITIVE,,,

SO,WE CAN SAY THAT,,,,,, LIM X TENDS TO INFINITIVEF(X)=LIM X TENDS TO INFINITIVE F(X+1)=LIM X TENDS TO INFINITIVE F(X+2)=T

SO,THAT

3T=T+5/T;

2T=5/T;

THEREFORE;T=UNDER ROOT 5/2

THUS.T=F(X)=UNDER ROOT 5/2,,,,

I HOPE YOU CAN UNDERSTAND,,,,,,,,,